OpenStudy (anonymous):

y^2 = 16x Write the equation using polar coordinates (r, theta).

OpenStudy (anonymous):

r^2sin^2t=16r cos t devide everything by r sin^2t: r=16cotan t cosec t

OpenStudy (anonymous):

So is that in the form (r, theta)?

OpenStudy (anonymous):

If it wasn't clear, @myko used the substitution \(y=r\sin t\) and \(x=r\cos t\).

OpenStudy (anonymous):

Okay, that makes sense

OpenStudy (anonymous):

Thanks